OpenStudy (anonymous):

Jake has a bag of 50 beads, some of which are blue and the remaining are green. Jake randomly pulls out a bead from the bag, records the color, and replaces it in the bag. Jake has already recorded 6 blue and 14 green beads. Based on these results, what is most likely the number of blue beads in the bag?

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

what are the options

OpenStudy (anonymous):

its a writing question

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

well what do you think?

OpenStudy (igreen):

Well, he drew 20 times.

OpenStudy (igreen):

I guess we could set up a proportion..

OpenStudy (igreen):

\(\sf\dfrac{6}{20}=\dfrac{x}{50}\) Because on 20 tries we got 6 blue beads and we'll find out how much we get with 50 blue beads(which is how many we have.

OpenStudy (igreen):

Multiply 50 to both sides: \(\sf x = \dfrac{6\times50}{20}\) Simplify

OpenStudy (anonymous):

15% @iGreen